OpenStudy (kylaaray2004yay):

A miner descended 1200 feet into a mine. While in the mine, he then rose 900 feet and then descended 350 feet. What was the miner's net gain or loss in altitude? A. –650 ft B. 50 ft C. 650 ft D. –1200 ft

OpenStudy (hafsahh):

A

OpenStudy (zzr0ck3r):

Don't just give answers @hafsahh explain and guide.

OpenStudy (hafsahh):

descend means to go down therefore the value 1200 is actually -1200 -1200 +900-350 -1200+900=-300 -300-350=-650